the paint in a certain container is sufficient to paint an area equal to 9.375 meter square.how many bricks of dimensions 22.5cm*10cm*7.5cm can be painted out of this container?
Answer by 303795(602)

To solve this problem find the surface area of a brick in square metres then divide the area covered by the paint by the surface area of one brick.
The brick is (0.225m x 0.10m x 0.075m )
Surface area of the top (and bottom) is 0.225*0.10=0.0225 square metres.
Area of the front (and rear) is 0.225 * 0.075 = 0.016875 Square metres.
Areas of the end (same for other end) is 0.10 * 0.075 = 0.0075 square metres
Total surface area of a brick is 2*0.0225 + 2*0.016875 + 2*0.0075 = 0.09375 square metres
Bricks to be painted = 9.375 /0.09375 = 100